OpenStudy (anonymous):

Find the equation of the tangent lone to the curve y=x^3+4x^2+4 at (1,9)

OpenStudy (accessdenied):

1) Find the derivative of y with respect to x 2) Plug in the x-value of point into derivative equation for slope of line at x 3) Use point-slope form to find line equation y - y1 = m(x - x1)

OpenStudy (rogue):

You can use the point slope form, or you can use the linearization form, which I find more efficient. The tangent line to the curve f(x) at (Xo, Yo is\[Y_{tangent} = f'(x_0)(x - x_0 )+y_0\]

OpenStudy (anonymous):

would the answer be y=11x-2

OpenStudy (rogue):

Yup, good job! =)

OpenStudy (anonymous):

would you be able to show me how to do it step by step?

OpenStudy (rogue):

Yeah, sure, hold on.

OpenStudy (rogue):

Tangent line to the curve y=x^3+4x^2+4 at (1,9). Step 1: Find y'(1)\[ y=x^3+4x^2+4\]\[y' = \frac {d}{dx} \left[ x^3+4x^2+4 \right]\]\[y' = 3x^2 + 8x\]\[y' (1) = 3 * 1^2 + 8 *1 = 11\] Step 2: Plug in your info. into tangent line equation. You have f'(1) = 11, the x and y coordinates the line passes to be x = 1, and y = 9.\[Y_{tangent} = f'(x_0)(x - x_0 )+y_0\]\[Y_{tangent} = 11*(x - 1 )+9 = 11x - 11 + 9 = 11x - 2\] Done! Does that help?

OpenStudy (anonymous):

thank you i did it correct really appreciate it (:

OpenStudy (rogue):

Alrighty, awesome :)